# discrete-time diffusion-like models for speech synthesis

### v1

@@ 4 Experiments
Our work follows the pipeline of Grad-TTS, which includes an encoder $E_{\theta}(\mathbf{c})$, a decoder $R_{\theta}(.,E_{\theta}(\mathbf{c}))$, and a fixed HIFI-GAN [30] vocoder. Each of our systems is trained by using an encoder $E_{\theta}(\mathbf{c})$ checkpoint from Grad-TTS, and training the decoder from scratch for 500 epochs with a frozen encoder. Then, the encoder is unfrozen and trained for another 100 epochs. The baseline model uses a checkpoint provided by Grad-TTS with either 5 or 10 inference steps[7]. Specifically, Grad-TTS is a score-based model which learns scores and integrates the score over a discretized SDE for inference. On the other hand, our implemented models are non-score based, where clean data is predicted (Section 3.2) and used to refine the result (Algorithm 1, Algorithm  2) in each iterative step during inference (Section 3.3. The training and inference conditions are consistent in our models, meaning there is no discretisation error.
@@ 4.2 Subjective evaluation
The subjective evaluation is carried out on the Amazon Mechanical Turk (AMT) platform. 30 randomly selected texts from test set are used throughout the subjective evaluation. Each system is evaluated on 30 distinct speech samples, each rated by 19 native speakers. The ground_truth result is generated by the fixed HIFI-GAN [30] vocoder.
@@
Ground_truth $4.22$
Baseline $3.89$
Grad-TTS-DT $3.87$
RFAG $(\sigma=0.6)$ $3.92$
RFMG $(\sigma=0.6)$ $3.88$
Blurring $3.87$
Mixture $3.91$
@@
All our systems achieved similar MOS score to the baseline. It is also shown by Fig. 1 that the MOS results are distributed similarly across the different MOS levels in all systems. In MOS evaluation, the RFAG and Mixture slightly outperformed the baseline, and it shows the effectiveness of discrete time diffusion-like models. Also, the MOS of RFMG is only 0.01 below the baseline, which shows multiplicative Gaussian noise can be a viable noising strategy. In addition, the encoder used in our systems is suboptimal and we believe that with tuning most of our systems could perform better.

### v2

@@ 4 Experiments
Our work follows the pipeline of Grad-TTS, which includes an encoder $E_{\theta}(\mathbf{c})$, a decoder $R_{\theta}(.,E_{\theta}(\mathbf{c}))$, and a fixed HIFI-GAN [30] vocoder. Each of our systems is trained by using an encoder $E_{\theta}(\mathbf{c})$ checkpoint from Grad-TTS, and training the decoder from scratch for 500 epochs with a frozen encoder. Then, the encoder is unfrozen and trained for another 100 epochs. The baseline model uses a checkpoint provided by Grad-TTS with either 5 or 10 inference steps [7]. Specifically, Grad-TTS is a score-based model which learns scores and integrates the score over a discretized SDE for inference. On the other hand, our implemented models are non-score based, where clean data is predicted (Section 3.2) and used to refine the result (Algorithm 1, Algorithm  2) in each iterative step during inference (Section 3.3). The training and inference conditions are consistent in our models, meaning there is no discretization error.
@@ 4.2 Subjective evaluation
The subjective evaluation is carried out on the Amazon Mechanical Turk (AMT) platform. 30 randomly selected texts from test set are used throughout the subjective evaluation. Each system is evaluated on 30 distinct speech samples, each rated by 20 native speakers. The ground_truth result is generated by the fixed HIFI-GAN vocoder.
@@
Ground truth $4.15$
Baseline $4.02$
Grad-TTS-DT $4.07$
RFAG $(\sigma=0.6)$ $3.90$
RFMG $(\sigma=0.6)$ $3.89$
Blurring $3.86$
Mixture $3.86$
@@
All our systems achieved similar MOS scores to the baseline. It is also shown by Fig. 1 that, aside from Ground truth system, the MOS results are distributed similarly across the different MOS levels. In MOS evaluation, Grad-TTS-DT slightly outperformed the baseline, demonstrating the effectiveness of a fully discrete process compared with the partially discretized process. In addition, the encoder used in our systems is suboptimal and we believe that with tuning most of our systems could perform better.